# ♪  Something Just Like TRuST ♪ thanks: An allusion to Coldplay’s “Something Just Like This.” : Toxicity Recognition of Span and Target

### v1

@@ Abstract
Toxicity in online content, including content generated by language models, has become a critical concern due to its potential for negative psychological and social impact. This paper introduces TRuST, a comprehensive dataset designed to improve toxicity detection that merges existing datasets, and has labels for toxicity, target social group, and toxic spans. It includes a diverse range of target groups such as ethnicity, gender, religion, disability, and politics, with both human/machine-annotated and human/machine-generated data. We benchmark state-of-the-art large language models (LLMs) on toxicity detection, target group identification, and toxic span extraction. We find that fine-tuned models consistently outperform zero-shot and few-shot prompting, though performance remains low for certain social groups. Further, reasoning capabilities do not significantly improve performance, indicating that LLMs have weak social reasoning skills.
@@
Disclaimer: Due to the topic studied here, the paper contains offensive words.
@@ 1 Introduction
Offensive and toxic content is pervasive in social media and online forums. Because large language models (LLMs) are trained on online data, they learn to generate toxic output Gehman et al. (2020); Hartvigsen et al. (2022). Exposure to toxic content might lead to lack of empathy, prejudice Pluta et al. (2023), stress, and depression Saha et al. (2019). Social groups that might need special attention such as people with developmental disorders can be more vulnerable to toxic content Kim et al. (2023). Furthermore, LLMs are being used in high-stakes conversational applications such as education Yu et al. (2025); de Araujo et al. (2025) and healthcare Yeo et al. (2024, 2025). Exposure to offensive language can lead to issues with self-esteem, anger, and anxiety Kansok-Dusche et al. (2023). Within the NLP community there has been increasing interest in toxicity detection Davidson et al. (2017); Rawat et al. (2024); Duan et al. (2025); Khurana et al. (2025) and toxicity mitigation Suau et al. (2024); Ermis et al. (2024); Pozzobon et al. (2023a); Li et al. (2024); Liu et al. (2021). However, existing toxicity datasets differ in format, leading to lack of uniformity in benchmarking. We build on existing datasets to create TRuST to improve benchmarking, and to provide a more comprehensive picture of online use of toxic language.
Figure 1: Histogram of distinct tokens within toxic spans (x-axis) ordered by their total count (y-axis). Note that the figure has a vertical "tear" at around 150 on the x-axis to show continuation out beyond 500 distinct tokens, but with counts too low to be visible. The most common 25 words are shown descending on a diagonal.
Existing toxicity datasets for English differ in size, source, and annotation scheme. They can range from about 5K examples Zampieri et al. (2023) to four times that size Davidson et al. (2017); Mathew et al. (2021) for human-generated posts from social media (typically twitter), or an order of magnitude larger for AI-generated datasets Hartvigsen et al. (2022). Davidson et al. (2017) introduced a toxicity scheme for hate speech, offensive language, or neither that has been widely adopted Almohaimeed et al. (2023); Mathew et al. (2021), although binary labels for offensiveness have also been used and combined with hierarchical schemes for the target of offense Zampieri et al. (2019). Categorizations of target include higher level classes for individual, group or other Zampieri et al. (2019), as well as very fine-grained labels, thus the THOS dataset Almohaimeed et al. (2023) has 31 targets (called topics). Some datasets also include identification of the spans of words that support identification of toxicity or offensiveness within a post or comment Zampieri et al. (2023); Mathew et al. (2021); Pavlopoulos et al. (2022). We review existing datasets and their annotation schemes to arrive at a synthesis that includes toxicity, target and span. In particular, we define a closed set of social group targets that includes half a dozen higher level classes and over 20 fine-grained subgroups. Our efforts produce TRuST, a large scale dataset of over 300K examples, created by merging many existing datasets that represent diverse social media sources as well as AI-generated examples. We also train human annotators to label a subset of 11,500 examples, reaching interannotator agreement scores in line with previous work.
To investigate the benefits of TRuST, we carry out three activities: an analysis of the characteristics of toxic language in social media (cf. Fig. 1); benchmark tests of LLMs on toxicity detection; automatic annotation of all the TRuST examples apart from the manually annotated subset, using the best method from our benchmarking tests. We benchmark LLMs using zero-shot, few-shot, and Chain-of-Thought prompts and compare LLMs to pre-trained language models (PLM) that we fine-tune. We find that fine-tuned PLMs perform somewhat better than LLMs. We also find that performance remains low on certain social groups, suggesting that social reasoning in LLMs needs improvement.
Our contributions are: (1) TRuST, a comprehensive dataset with toxicity, target social group, and toxic span labels; (2) a comprehensive definition of toxicity based on a synthesis of previous work; (3) benchmarking of SOTA LLMs, and their comparison with previous SOTA models.
@@ 2 Related Work
2.1 Definitions of Toxicity
Since the perception of toxicity depends on factors such as context, background, demographics, it is subjective. This has led to work that provides parallel annotations of toxicity from different demographic groups Mostafazadeh Davani et al. (2024). Further, types of toxicity annotation have ranged from narrower categories such as hate speech to broader types of offense.
Davidson et al. (2017) define hate speech HS as hatred towards a social group or humiliation/insult to the members of a group. Additionally, HS might include threats or reference to violence. Nockleby (2000) defines HS as any communication that disparages a person or a group on the basis of a social characteristic such as race, color, ethnicity, gender, sexual orientation, nationality, religion, or other characteristic. These works generally agree that HS is directed at a target.
Other types of toxic content go beyond hate speech. Dorn et al. (2024) state that toxic content includes using slurs in a pejorative way, attacking or criticizing a minority, promoting hate speech, negative stereotypes, etc. The definition used in PerspectiveAPI Lees et al. (2022) for toxic content is “rude, disrespectful or unreasonable language that is likely to make someone leave a discussion”. Kumar et al. (2021) propose that toxicity involves identity-based attacks such as racism, bullying, threats of violence, sexual harassment, and includes emotional harm and hate speech.
One issue in the literature about toxicity is that some terms are used interchangeably, or one of them is seen as a subcategory of the other. Therefore, there is a lack of consensus which leads to confusion and inconsistency Khurana et al. (2022). Moreover, some works consider every content with slurs or swear words as toxic, which may not be the case (e.g. you are f***ing beautiful).
We take toxicity to comprise three categories: hate speech, abusive language, and sexual harassment. Hate speech is defined as offensive and discriminatory discourse towards a group or an individual based on characteristics such as race or religion, thus always has a target. It includes negative stereotyping (negative traits and characteristics attributed to a group), racism (discrimination against or negative attitudes towards individuals/groups belonging to a certain race), sexist language (discrimination or behaviors that foster stereotypes based on a gender), and discrimination based on sexual orientation. Abusive language is content with inappropriate words such as profanity or disrespectful terms for people based on sociodemographic characteristics. It includes psychological threats (expressions of an intent of harms such as humiliation, causing distress, constant criticism). Our last category is sexual harrasment which includes unwelcome sexual moves, requests of sexual favors, or other unwanted physical/verbal behaviors of a sexual nature towards someone. In our work, toxic language often has a target, but can also involve use of offensive words in an aggressive fashion without targeting a specific social group, e.g., “honestly? I can handle kpop stans dragging armys but just stay the f**k away from bts they’ve done lit rally nothing to y’all.” Our annotation instructions (see Appendix C) include a binary label for toxicity that covers these three categories.
2.2 Datasets
Davidson et al. (2017) produced one of the earliest datasets focusing on hate speech and offensiveness in social media. Their dataset has around 25k tweets with three-way labels for hate speech, offensive language or neither, assigned by crowdworkers. OLID Zampieri et al. (2019) has 14k tweets and hierarchical offensiveness labels. The first level is offensive or not. Offensive examples are then labeled as targeting a group or not. The last level is for the targeted offensive tweets for whether the target is an individual, social group, or any other category such as an organization.
Inclusion of annotations for target social group supports deeper understanding of which are the groups that are targeted most often, and other bias-related issues. TBO Zampieri et al. (2023) is a recent dataset that has target and toxic span annotations, both of which support explainability. However, their target annotation involves identifying specific words in toxic language samples, rather than more general categories, which limits their utility. The THOS hate speech dataset Almohaimeed et al. (2023) has two-level categories for target: the first is more high level such as country or religion (6 classes), and the second is more detailed belonging to the first level categories (31 classes) such as a term for a country (China), practitioner of a religion (Muslim), or racial group (Black). Similar to THOS, we employ two levels with 8 higher levels (including No Target), 5 of which are broken down into two or more subclasses.
Another annotation type that supports explainability is the span of words that constituted the offense, as in ToxicSpans Pavlopoulos et al. (2022). HateXplain Mathew et al. (2021) has target group annotations based on religion, gender etc. and rationales that are the spans on which annotators’ labeling decision is based. The rational annotation can contain words for target group. Our set of targets is more diverse in the number of target groups.
The cited works mostly focus on explicit hate speech; implicit hate speech is harder to detect. ElSherief et al. (2021) propose a new dataset with 19k tweets with an implicit versions of hate and toxicity (e.g., irony). Social Bias Frames Sap et al. (2020) creates a formalism that models pragmatic frames. They have annotations for targeted group, implied statement, offensiveness, use of in-group language etc. Alongside datasets with social media texts, some works create data using generative language models, such as ToxiGen Hartvigsen et al. (2022). The latter focuses on implicit hate speech; they generate both toxic and benign examples for 13 minority groups using GPT-3 Brown et al. (2020). Their dataset has 274k examples and labels are based on the prompts used for GPT-3 with the assumption that if the prompt is toxic, the generated text is also toxic. FairPrism Fleisig et al. (2023b) is another machine-generated dataset that builds on ToxiGen and Social Bias Frames. They collect human annotations on stereotyping and demeaning harms, target social group, and providing non-aggregated annotations to support fairness.
Some argue that the context, author, and reader are important factors for perception of toxic language Cowan and Hodge (1996); Nieto and Boyer (2006). This motivates Zhou et al. (2023) to use GPT3.5 to build CobraFrames, with contextual annotations for speaker and listener, target group, and emotional reactions or power dynamics between the speaker and listener.
@@ 2.3 Prediction of Toxicity, Targets and Spans
The SOTA for toxicity detection is PerspectiveAPI Lees et al. (2022), in which multilingual BERT Devlin et al. (2019) is first trained, then multiple single-language convolutional neural networks are distilled. However, it has been criticized for having only moderate correlation with humans Welbl et al. (2021); Schick et al. (2021), and over-reliance on surface patterns such as swear words, resulting in a high rate of false positives Rosenblatt et al. (2022). Another issue is that they sometimes release updates without notification or explanation of improvements, making previous findings questionable Pozzobon et al. (2023b).
A common approach for target social group is a multi-layer perceptron classifier whose input is embeddings of examples from a pretrained encoder, such as BERT Mathew et al. (2021). LLMs have also been used Zhou et al. (2023).
@@
Table 1: Statistics for our human annotated data showing the total count (and percentage of the total) for each higher level or lower-level social group, and the percentage of each that are labeled toxic. Lower-level groups with the highest and lowest proportion of toxic texts are in red and green font, respectively. The last two columns are for the test set (T.). In targets, o. means other, native means native american.
Toxic span prediction is usually treated as a multi-token classification problem where a label is predicted for each token He et al. (2024). SpanBERT Joshi et al. (2020), a pretrained model for high quality embeddings for predicting spans of text, is the SOTA encoder for this task. While there is little work on using LLMs for toxic span extraction, it has been done for Romanian toxic span identification using GPT4 Paraschiv et al. (2023).
3 Dataset
TRuST re-annotates data from ToxicSpan (CC-BY 4.0) Pavlopoulos et al. (2022), HateXplain (CC-BY 4.0)Mathew et al. (2021), TBO (CC-BY 4.0) Zampieri et al. (2023), Thos (CC-BY 4.0) Almohaimeed et al. (2023), and ToxiGen (CC-BY 4.0) Hartvigsen et al. (2022). We observed some noise in the annotations, which is inevitable due to the subjective nature of toxicity. The ToxiGen toxicity labels, however, were particularly noisy, due to the strong assumption of assigning the prompt labels to the generated text. Where necessary, we added annotations for target group and span. TRuST therefore has three types of labels, toxicity, target group and span, throughout. The full dataset has nearly 300K examples (see Table A).
To provide a high-quality benchmark for testing toxicity models, we collected human annotations for binary toxicity, target social group, and toxic spans for a subset of over 10k examples. We have 24 target social groups, including “no target”. Note that we separate “Chinese” from Asian because it includes a significantly large number of examples. After testing different models on prediction of the three types of labels, we annotate the rest of the data with the best-performing models.
3.1 Human Annotation Procedure
To collect the manual annotations, we hired six undergraduate students studying computer science or data science who have some experience with data analysis; they were paid \$$\$10$/hour. We recruited students from different backgrounds (Indian, Chinese, White) and demographics to have more diversity, given the cultural and demographic differences in toxicity Mostafazadeh Davani et al. (2024); Fleisig et al. (2023a). They were provided with detailed instructions (see Appendix C). To ensure good quality, we did three iterations of annotator training. We should note that by the time we collected toxic span annotation, which was done last, only three of the original six annotators were still available. Our inter-annotator agreement scores using Krippendorf’s $\alpha$ Krippendorff (2013) are 0.56, 0.66, and 0.55 for toxicity, target social group and spans, respectively. Agreement scores in previous work (Krippendorff’s $\alpha$ unless otherwise noted) are, for toxicity: 0.46 Mathew et al. (2021), 0.51 Sap et al. (2020) and 0.64 Hartvigsen et al. (2022); for target social group: 0.50 Sap et al. (2020); for spans: 0.55 Cohen’s kappa Pavlopoulos et al. (2022)). To calculate agreement for spans, we incorporate MASI Passonneau (2006), a distance metric for inter-annotator agreement on sets, based on a weighted Jaccard, where the weight is higher for set subsumption than for set intersection, which in turn is higher than for disjunction.
After annotators were trained, they first annotated target social group for every example (including no target), followed by toxicity. During the toxicity annotation, they also double-checked the previously labeled target social group. The examples where a previous target social group annotation was considered incorrect were labeled by all students, where we applied majority voting to get the final label. Span annotation came last, during which students did a sanity check on the toxicity label, where those judged incorrect were again relabeled by the three remaining annotators, with majority voting for the final label. Notably, span annotation identified either individual words that led to the judgment that the example was toxic, or applied to the entire sentence.
3.2 Toxicity and Target Group Labels

### v2

# ♪ Something Just Like TRuST ♪ An allusion to Coldplay’s ‘‘Something Just Like This.’’ : Toxicity Recognition of Span and Target

@@ Abstract
Toxic language includes content that is offensive, abusive, or that promotes harm. Progress in preventing toxic output from large language models (LLMs) is hampered by inconsistent definitions of toxicity. We introduce TRuST, a large-scale dataset that unifies and expands prior resources through a carefully synthesized definition of toxicity, and corresponding annotation scheme. It consists of $\sim$300k annotations, with high-quality human annotation on $\sim$11k. To ensure high-quality, we designed a rigorous, multi-stage human annotation process, and evaluated the diversity of the annotators. Then we benchmarked state-of-the-art LLMs and pre-trained models on three tasks: toxicity detection, identification of the target group, and of toxic words. Our results indicate that fine-tuned PLMs outperform LLMs on the three tasks, and that current reasoning models do not reliably improve performance. TRuST constitutes one of the most comprehensive resources for evaluating and mitigating LLM toxicity, and other research in socially-aware and safer language technologies.
@@
Disclaimer: Due to the topic studied here, the paper contains examples of offensive language.
@@
Figure 1: Histogram of distinct tokens within toxic spans (x-axis) ordered by total count (y-axis). The vertical "tear" at around 150 on the x-axis shows there is a long tail. The 25 most common words are shown descending on a diagonal.
Dataset Size Text Source Tox. High-level Target Fine-grained Target Span
fleisig-etal-2023-fairprism 5k LLM 3 2 Not Provided
zampieri-etal-2019-predicting 14k Social Media 2 3
elsherief-etal-2021-latent 19k Social Media 8 free text
davidson2017automated 25k Social Media 3
zhou-etal-2023-cobra 33k LLM free text free text
sap-etal-2020-social 45k Social Media 4 free text
zampieri-etal-2023-target * 4.5k Social Media 2 free text
almohaimeed2023thos * 8.3k Social Media 4 6 31
mathew2021hatexplain * 20k Social Media 3 5 18
pavlopoulos-etal-2022-detection * 11k Social Media 1
hartvigsen-etal-2022-toxigen * 274k LLM 2 13
[…]
Table 1: Comparison of datasets by size, source, and annotation scheme. Social Media includes Twitter, Gap, and Civil comments; LLMs were GPT3 and GPT3.5.
@@ 1 Introduction
Toxic and offensive language is pervasive online, and because LLMs are trained on web data, they sometimes generate such content gehman-etal-2020-realtoxicityprompts; hartvigsen-etal-2022-toxigen. The resulting exposure of users to toxic output can reduce empathy toward targeted people, reinforce social biases, and inflict direct harms on targeted communities. Identity-directed toxicity can reinforce discriminatory attitudes pluta2023exposure, while repeated exposure of vulnerable users can cause psychological burdens, such as stress or depression saha2019prevalence, and can contribute to anger issues kansok2023systematic. As LLMs enter high-stakes interactive settings such as education deAraujoEtAl25, harmful outputs (e.g., abusive tutoring responses such as “lol only a female n*** could be so dumb”) can propagate these harms to children. Accordingly, many mitigation methods have been proposed suau2024whispering; ermis-etal-2024-one; pozzobon-etal-2023-goodtriever; li-etal-2024-preference; liu-etal-2021-dexperts. Research on toxicity detection, however, has been evaluated on datasets that differ widely in scale, source, and annotation schemes davidson2017automated; rawat2024hate; duan-etal-2025-exploring; khurana-etal-2025-defverify. This limits comparability and hampers improvement. Progress depends on reliable toxicity measurement, which motivates our work on a rigorous merger of existing datasets combined with comprehensive benchmarking.
Table 1 illustrates the diversity in size and source of existing datasets, ranging from $\sim$5K examples to much larger collections of human-authored social media versus LLM-generated data. Annotation labels include binary toxicity to multi-class taxonomies (e.g., profanity vs. hate speech), and differ in whether annotation elements include the target social group or toxic words. Even the definitions of “toxicity” differ: pavlopoulos-etal-2022-detection includes offensive language, while almohaimeed2023thos does not. A more unified annotation scheme that captures key components of toxicity could support comparability of results and overcome observed flipping of model rankings across studies (e.g., trager-etal-2025-mftcxplain vs. singh2025rethinking). TRuST balances scale and label quality, and adopts a unified annotation framework.
We curate the five datasets shown with asterisks in Table 1, and merge them as described in a later section. TRuST contains $\sim$300K examples with binary toxicity, target social group, and toxic span annotations across diverse sources, plus a human-annotated subset. Although our primary annotators were college students with similar educational backgrounds and limited gender diversity, we validated annotation quality using a diverse volunteer group: agreement is good within the primary group (Krippendorf $\alpha$ 0.56/0.66/0.55 for toxicity/target/span), within volunteers (0.61/0.62/0.47), and across groups (0.63/0.62/0.45). These reliabilities match or exceed prior reports (0.46 toxicity mathew2021hatexplain, 0.50 target sap-etal-2020-social, 0.55 spans pavlopoulos-etal-2022-detection).
To investigate TRuST’s value, we analyzed toxic language patterns in the social media data. Figure 1 shows that context-dependent terms (e.g., ”hate” or ”sand”) appear frequently in toxic text, where surrounding context determines whether usage is toxic (e.g., “UUGHHHH typical arab shit man !!! I hate those sand people”). At the same time, the distribution has a long tail: most toxic terms occur only a few times. This highlights that toxicity detection requires nuanced, context-aware modeling and that toxic span prediction is inherently challenging.
Using TRuST, we benchmarked SOTA methods including PerspectiveAPI, ShieldGemma, LlamaGuard, LLMs, and fine-tuned PLMs in toxicity classification, target group prediction and toxic span identification. For LLMs, we used different prompting strategies including zero/few-shot and Chain-of-Thought. We found that PerspectiveAPI, ShieldGemma, and LlamaGuard do not generalize well to our dataset, and that fine-tuned PLMs outperform all methods in all three tasks. All methods have toxicity detection performance differences, depending on the target group. Further, reasoning methods, including CoT and reasoning-enhanced models, do not help, indicating the need for improvements in social reasoning. Interestingly, models performed better on LLM-generated data, indicating that social media data is more diverse and challenging.
Lastly, we utilized the best method from our benchmarking tests to increase the size of TRuST with synthetic labels. We show that models trained on TRuST generalize well to OpenAI dataset markov2023holistic, which is an independent dataset with human and machine-generated text.
In sum, our three contributions are: (1) a comprehensive definition of toxicity and an annotation scheme synthesized from previous work; (2) the TRuST dataset labeled for toxicity, target group, and toxic span; (3) extensive benchmarking across state-of-the-art toxicity detection methods.
@@ 2 Related Work
In this section, we discuss differences in definitions for and annotation of toxicity. In addition, we review state-of-the-art methods for prediction of toxicity, target group, and toxic span.
2.1 Annotation of Toxicity
Prior work on toxic language annotation adopts different definitions and criteria for what counts as “toxic” or “hate” speech, which leads to comparability issues across datasets and models khurana-etal-2022-hate. To reduce subjectivity, other elements are often introduced, such as (i) the target group and (ii) the words that justify the toxicity decision (i.e., rationales, or spans). Works differ regarding which behaviors count as toxic, e.g., threats, humiliation, stereotyping) davidson2017automated; john2000hateversus whether it includes rude, disrespectful or abusive language, and identity-based attacks lees2022newgenerationperspectiveapi; kumar2021designing; dorn2024harmful. Here, we review sources of inconsistency in toxicity, target, and span annotation to motivate our design choices.
Datasets differ also in the granularity of toxicity labels. It can be binary, as in pavlopoulos-etal-2022-detection; zampieri-etal-2023-target; zampieri-etal-2019-predicting; hartvigsen-etal-2022-toxigen, or a multiclass label davidson2017automated; mathew2021hatexplain; almohaimeed2023thos; sap-etal-2020-social; elsherief-etal-2021-latent. In TRuST, we adopt a binary toxicity label with clear guidelines grounded in three broad categories: hate speech, abusive language, and sexual harassment.
john2000hate and davidson2017automated agree that hate speech is directed at a target, yet explicit target labels are rarely given. As illustrated in Table 1, almohaimeed2023thos includes 31 fine-grained targets, whereas others annotate arbitrary free text sap-etal-2020-social, or omit targets entirely. In TRuST, we introduce 8 high-level target groups with 24 fine-grained subgroups, to support standardized analysis and more consistent reporting across models.
Finally, only a small subset of existing works, such as pavlopoulos-etal-2022-detection; mathew2021hatexplain; zampieri-etal-2023-target, include span-level labels to explain toxicity. Even among these, there is variation in annotation of toxic trigger words vs. annotator-provided rationales. The absence of spans encourages models to rely on surface cues (e.g., swear words) and reduces explainability. In TRuST, we provide token-level toxic span annotation, allowing spans to cover either specific triggers or the whole sentence (e.g., for implicit toxicity). This supports explainable evaluation and helps reduce over-reliance on isolated slur words. Our span annotation of whole sentences is novel and helps identify more nuanced cases.
2.2 Dataset Sources
Human-labeled and generated datasets are usually higher quality but more costly and time-consuming than machine-generated data. However, machine-generated data is usually less diverse, and labels are noisy when machine-generated. As you see in Table 1, previous datasets are mostly human-generated and labeled, but the annotation schemes differ. We aim to combine the strengths of both approaches by establishing a consistent annotation scheme for merging and re-annotating data drawn from previous datasets, with careful inter-annotator reliability and impact of diversity.
@@ 2.3 Prediction of Toxicity, Targets and Spans
Comparison of prior toxicity prediction methods is challenging because datasets use highly variable annotation schemes khurana-etal-2022-hate (see Table 1). For toxicity detection, Perspective API lees2022newgenerationperspectiveapi has long been considered SOTA due to its accessibility and multilingual availability, producing probability scores for categories of offensive language (e.g., toxicity, insult). However, it shows only moderate correlation with human judgments welbl-etal-2021-challenges-detoxifying; schick2021self and over-relies on surface cues such as swear words, leading to many false positives rosenblatt2022critical. Further, LLM-based guards such as ShieldGemma zeng2024shieldgemma and LlamaGuard inan2023llama are trained to classify prompts and/or responses into safety-policy categories (or binary allow/block decisions). Their evaluations are typically reported on moderation benchmarks (e.g., OpenAI Moderation markov2023holistic, ToxicChat lin2023toxicchat. Fine-tuned PLMs (e.g., BERT/RoBERTa) can report high toxicity accuracy, up to 0.95 on some datasets tsai2025fine; joshi2025enhancing; however, the best PLM accuracy is notably lower (0.79) and PerspectiveAPI performs near chance (0.50) (cf. Table 3) on TRuST, highlighting sensitivity of “SOTA” claims to dataset.
Target-group prediction is most commonly modeled by training a lightweight classifier (e.g., linear/MLP head) on top of pretrained encoders such as BERT/RoBERTa mathew2021hatexplain. LLMs have also been used for more context-dependent target/explanation generation zhou-etal-2023-cobra.
Finally, toxic span prediction is typically framed as token-level classification he2024you, where SpanBERT is a strong encoder due to its span-oriented pretraining joshi2020spanbert. Prior LLM-based span results are limited (e.g., GPT-4 for Romanian with F1 0.72)
Overall, while toxicity detection and moderation have advanced rapidly, the lack of a consistent annotation scheme and established benchmark across toxicity, target, and span makes it difficult to compare methods reliably; TRuST is designed to correct this lack of reliable toxicity measurement.
3 Dataset Assembly
TRuST draws from the datasets in the last 5 rows of Table 1, based on the criteria of datasets wtih span annotation, or with multiple target groups. We kept all text from these datasets, and re-annotated it with our framework. ToxiGen labels were particularly noisy, due to assigning the prompt labels to the generated text; Krippendorf’s $\alpha$ was only 0.37 between our human annotation and ToxiGen labels. Here, we present our definition of toxicity, describe our human annotation, and characterize the dataset.
3.1 Definition of Toxicity
Based on our review in the previous section, we take toxicity to comprise three broad categories: hate speech, abusive language, and sexual harassment. Hate speech is defined as offensive and discriminatory discourse towards a group or individual based on characteristics such as race or religion, thus always has a target. It includes negative stereotyping (negative traits attributed to a group), racism (discriminatory actions or attitudes towards based on race), sexist language (fostering stereotypes based on a gender), and discrimination based on sexual orientation.
Abusive language is content with inappropriate words such as profanity or disrespectful terms for people based on sociodemographic characteristics. It includes psychological threats, meaning expressions of an harms such as humiliation, intent to cause distress, or criticism motivated by bias.
Our last category is sexual harrasment which includes unwelcome sexual moves, requests of sexual favors, or other unwanted physical/verbal behaviors of a sexual nature. In our work, toxic language often has a target, but can also involve use of offensive words in an aggressive fashion without targeting a specific social group, e.g., “honestly? I can handle kpop stans dragging armys but just stay the f**k away from bts they’ve done lit rally nothing to y’all.” Our annotation instructions (see Appendix C) include a binary label for toxicity defined in terms of these three categories.
3.2 Human Annotation Procedure
We collected human annotations for binary toxicity, target social group, and toxic spans for a subset of $\sim$11K examples. Although 87% of our dataset consists of LLM-generated text, to have more balance, our human annotated subset has only 40% LLM-generated text. Our primary annotator pool was only moderately diverse, therefore, to assess the impact of annotator diversity on reliability, we annotated a smaller subset ($N=300$) with a more diverse volunteer pool. We defined 24 target groups (including no target) and separated Chinese from Asian due to its high frequency.
We hired six undergraduate CS/data-science students (paid $\mathdollar 10$/hour) with data-analysis experience. Given reports of demographic differences in toxicity perception mostafazadeh-davani-etal-2024-d3code; fleisig-etal-2023-majority, we aimed for as much ethnic diversity as possible, ending with Indian, Chinese, and American/European. Annotators followed detailed guidelines emphasizing context over words alone (cf. Appendix C). They completed three training iterations. Due to dropout, span annotation was done by three annotators.
@@ 3.2 Human Annotation Procedure
Table 2: Statistics for our human annotated data showing the total count (and percentage of the total) for each higher level or lower-level social group, and the percentage of each that are labeled toxic. Lower-level groups with the highest and lowest proportion of toxic texts are in red and green font, respectively. The last two columns are for the test set (T.). In targets, "o." means other, "native" means native american.
After training, annotators labeled targets first, then toxicity (which can depend on the target), and finally spans, with consistency checks throughout. Following fleisig-etal-2023-majority, we instructed annotators to judge toxicity from the target group’s perspective; they also verified the target label during toxicity annotation. For spans, annotators marked the words justifying toxicity or, when toxicity was implicit (e.g., idioms, sarcasm, euphemisms), selected the full sentence elsherief-etal-2021-latent; wen2023unveiling; kim-etal-2024-lifetox. During span annotation, toxicity labels were checked and disagreements were relabeled and resolved via majority vote.
Krippendorff’s $\alpha$ was 0.56/0.66/0.55 for toxicity/target/span, respectively. For spans, we incorporated the MASI distance metric passonneau2006measuring, a weighted Jaccard with distinct weights for set subsumption >> intersection >> disjunction. This compares well with prior inter-annotator agreement measures of toxicity: 0.46 mathew2021hatexplain, 0.51 sap-etal-2020-social, 0.64 hartvigsen-etal-2022-toxigen; of target: 0.50 sap-etal-2020-social; of spans: 0.55 Cohen’s $\kappa$ pavlopoulos-etal-2022-detection).
To assess the effect of annotator diversity on reliability, we recruited 8 volunteers with diverse ethnicity (Caucasian, Arab, Asian), ages (18–32), and backgrounds (e.g., Nursing, Linguistics, Biobehavioral Health, CS), spanning multiple sexual orientations (e.g. homosexual, bisexual, and pansexual). They relabeled 300 randomly selected examples, yielding agreement of 0.61/0.62/0.47 for toxicity/target/span. Compared to agreement scores listed above, toxicity agreement is better but the other tasks are worse. Including the original labels gives similar agreement (0.63/0.62/0.45), suggesting our primary annotator pool is sufficiently reliable for the 11K human-annotated subset.
3.3 Human Annotated Subset